A new machine is expected to cost $6000 and have a life of 5 years . maintenance cost will be $1500 the first year , $ 1700 the second year , $ 1900 the third year , $2200 the forth year , and $2300 the fifth year . how much should be deposited in a fund that earns 9% per year , compounded monthly , in order to pay for this machine ?
1
Expert's answer
2020-04-27T07:43:36-0400

Year 0:



"\dfrac{\$6000}{\left(1 + \dfrac{0.09}{12}\right)^{0\times 12}} = \$6,000"

Year 1:

"\dfrac{\$1500}{\left(1 + \dfrac{0.09}{12}\right)^{1\times 12}} = \$1,371.34"

Year 2:

"\dfrac{\$1,700}{\left(1 + \dfrac{0.09}{12}\right)^{2\times 12}} = \$1,420.91"

Year 3:



"\dfrac{\$1900}{\left(1 + \dfrac{0.09}{12}\right)^{3\times 12}} = \$1,451.88"

Year 4:



"\dfrac{\$2200}{\left(1 + \dfrac{0.09}{12}\right)^{4\times 12}} = \$1,536.95"


Year 5:



"\dfrac{\$2300}{\left(1 + \dfrac{0.09}{12}\right)^{5\times 12}} = \$1,469.01"


Therefore, the total amount that should be deposited in a fund that earns 9% per year , compounded monthly , in order to pay for this machine is:


$6,000 + $1,371.34 + $1,420.91 + $1,451.88 + $1,536.95 + $1,469.01 = "\boxed{\color{red}{\$13,250.09}}"